AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 305     Entry time: Sun Feb 27 13:34:24 2022
Author: RSS TD 
Subject: Current status 

14:33: 95K-->140K

14:44: 140K --> 95K

Ring current ~6e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.43 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

ELOG V3.1.4-unknown